# Offline Safe Reinforcement Learning Using Trajectory Classification

### v4

@@ Appendix B Extended Results
BallRun ${\color[rgb]{.5,.5,.5}0.6\scriptstyle{\pm 0.1}}$ ${\color[rgb]{.5,.5,.5}5.08\scriptstyle{\pm 0.74}}$ ${\color[rgb]{.5,.5,.5}0.27\scriptstyle{\pm 0.14}}$ ${\color[rgb]{.5,.5,.5}1.46\scriptstyle{\pm 0.39}}$ ${\color[rgb]{.5,.5,.5}0.39\scriptstyle{\pm 0.09}}$ ${\color[rgb]{.5,.5,.5}1.16\scriptstyle{\pm 0.19}}$ ${\color[rgb]{.5,.5,.5}0.76\scriptstyle{\pm 0.01}}$ ${\color[rgb]{.5,.5,.5}3.91\scriptstyle{\pm 0.35}}$ ${\color[rgb]{.5,.5,.5}-0.47\scriptstyle{\pm 0.0}}$ ${\color[rgb]{.5,.5,.5}5.03\scriptstyle{\pm 0.0}}$ ${\color[rgb]{.5,.5,.5}0.22\scriptstyle{\pm 0.0}}$ ${\color[rgb]{.5,.5,.5}1.27\scriptstyle{\pm 0.12}}$ $0.59\scriptstyle{\pm 0.0}$ ${\color[rgb]{.5,.5,.5}3.52\scriptstyle{\pm 0.0}}$ ${\color[rgb]{0,0,1}\bm{0.27\scriptstyle{\pm 0.04}}}$ ${\color[rgb]{0,0,1}\bm{0.47\scriptstyle{\pm 0.19}}}$
CarRun $\bm{0.97\scriptstyle{\pm 0.02}}$ $\bm{0.33\scriptstyle{\pm 0.05}}$ $\bm{0.94\scriptstyle{\pm 0.0}}$ $\bm{0.22\scriptstyle{\pm 0.02}}$ ${\color[rgb]{0,0,1}\bm{0.99\scriptstyle{\pm 0.01}}}$ ${\color[rgb]{0,0,1}\bm{0.65\scriptstyle{\pm 0.31}}}$ $\bm{0.94\scriptstyle{\pm 0.01}}$ $\bm{0.15\scriptstyle{\pm 0.91}}$ ${\color[rgb]{.5,.5,.5}0.68\scriptstyle{\pm 0.01}}$ $7.78\scriptstyle{\pm 0.09}$ ${\color[rgb]{.5,.5,.5}0.95\scriptstyle{\pm 0.01}}$ $1.79\scriptstyle{\pm 0.18}$ $\bm{0.87\scriptstyle{\pm 0.0}}$ $\bm{0.0\scriptstyle{\pm 0.0}}$ $\bm{0.97\scriptstyle{\pm 0.0}}$ $\bm{0.03\scriptstyle{\pm 0.02}}$
DroneRun ${\color[rgb]{.5,.5,.5}0.24\scriptstyle{\pm 0.02}}$ ${\color[rgb]{.5,.5,.5}2.13\scriptstyle{\pm 0.62}}$ $\bm{0.28\scriptstyle{\pm 0.25}}$ $\bm{0.74\scriptstyle{\pm 0.97}}$ ${\color[rgb]{0,0,1}\bm{0.63\scriptstyle{\pm 0.04}}}$ ${\color[rgb]{0,0,1}\bm{0.79\scriptstyle{\pm 0.68}}}$ ${\color[rgb]{.5,.5,.5}0.72\scriptstyle{\pm 0.12}}$ ${\color[rgb]{.5,.5,.5}5.54\scriptstyle{\pm 0.81}}$ ${\color[rgb]{.5,.5,.5}0.42\scriptstyle{\pm 0.1}}$ ${\color[rgb]{.5,.5,.5}2.47\scriptstyle{\pm 0.34}}$ ${\color[rgb]{.5,.5,.5}0.33\scriptstyle{\pm 0.1}}$ ${\color[rgb]{.5,.5,.5}3.52\scriptstyle{\pm 0.58}}$ ${\color[rgb]{.5,.5,.5}0.67\scriptstyle{\pm 0.02}}$ ${\color[rgb]{.5,.5,.5}4.15\scriptstyle{\pm 0.1}}$ $\bm{0.55\scriptstyle{\pm 0.01}}$ $\bm{0.01\scriptstyle{\pm 0.0}}$
AntRun ${\color[rgb]{.5,.5,.5}0.7\scriptstyle{\pm 0.06}}$ ${\color[rgb]{.5,.5,.5}2.93\scriptstyle{\pm 2.4}}$ ${\color[rgb]{.5,.5,.5}0.65\scriptstyle{\pm 0.15}}$ ${\color[rgb]{.5,.5,.5}1.09\scriptstyle{\pm 0.84}}$ ${\color[rgb]{0,0,1}\bm{0.72\scriptstyle{\pm 0.04}}}$ ${\color[rgb]{0,0,1}\bm{0.91\scriptstyle{\pm 0.42}}}$ ${\color[rgb]{.5,.5,.5}0.76\scriptstyle{\pm 0.07}}$ ${\color[rgb]{.5,.5,.5}5.11\scriptstyle{\pm 2.39}}$ $\bm{0.15\scriptstyle{\pm 0.02}}$ $\bm{0.73\scriptstyle{\pm 0.07}}$ $\bm{0.03\scriptstyle{\pm 0.02}}$ $\bm{0.02\scriptstyle{\pm 0.09}}$ $\bm{0.61\scriptstyle{\pm 0.01}}$ $\bm{0.94\scriptstyle{\pm 0.69}}$ $\bm{0.67\scriptstyle{\pm 0.01}}$ $\bm{0.63\scriptstyle{\pm 0.1}}$
BallCircle ${\color[rgb]{.5,.5,.5}0.74\scriptstyle{\pm 0.15}}$ ${\color[rgb]{.5,.5,.5}4.71\scriptstyle{\pm 1.79}}$ $\bm{0.52\scriptstyle{\pm 0.08}}$ $\bm{0.65\scriptstyle{\pm 0.17}}$ ${\color[rgb]{.5,.5,.5}0.77\scriptstyle{\pm 0.06}}$ ${\color[rgb]{.5,.5,.5}1.07\scriptstyle{\pm 0.27}}$ ${\color[rgb]{.5,.5,.5}0.69\scriptstyle{\pm 0.11}}$ ${\color[rgb]{.5,.5,.5}2.36\scriptstyle{\pm 1.04}}$ ${\color[rgb]{.5,.5,.5}0.86\scriptstyle{\pm 0.18}}$ ${\color[rgb]{.5,.5,.5}3.09\scriptstyle{\pm 1.53}}$ ${\color[rgb]{.5,.5,.5}0.64\scriptstyle{\pm 0.01}}$ ${\color[rgb]{.5,.5,.5}0.76\scriptstyle{\pm 0.0}}$ ${\color[rgb]{.5,.5,.5}0.7\scriptstyle{\pm 0.04}}$ ${\color[rgb]{.5,.5,.5}2.61\scriptstyle{\pm 0.79}}$ ${\color[rgb]{0,0,1}\bm{0.68\scriptstyle{\pm 0.06}}}$ ${\color[rgb]{0,0,1}\bm{0.59\scriptstyle{\pm 0.07}}}$
CarCircle ${\color[rgb]{.5,.5,.5}0.58\scriptstyle{\pm 0.25}}$ ${\color[rgb]{.5,.5,.5}3.74\scriptstyle{\pm 2.2}}$ $\bm{0.5\scriptstyle{\pm 0.22}}$ $\bm{0.84\scriptstyle{\pm 0.67}}$ ${\color[rgb]{0,0,1}\bm{0.75\scriptstyle{\pm 0.06}}}$ ${\color[rgb]{0,0,1}\bm{0.95\scriptstyle{\pm 0.61}}}$ ${\color[rgb]{.5,.5,.5}0.63\scriptstyle{\pm 0.19}}$ ${\color[rgb]{.5,.5,.5}1.89\scriptstyle{\pm 1.37}}$ ${\color[rgb]{.5,.5,.5}0.74\scriptstyle{\pm 0.1}}$ ${\color[rgb]{.5,.5,.5}2.18\scriptstyle{\pm 1.33}}$ $\bm{0.71\scriptstyle{\pm 0.02}}$ $\bm{0.33\scriptstyle{\pm 0.0}}$ ${\color[rgb]{.5,.5,.5}0.49\scriptstyle{\pm 0.05}}$ ${\color[rgb]{.5,.5,.5}3.14\scriptstyle{\pm 2.98}}$ $\bm{0.64\scriptstyle{\pm 0.06}}$ $\bm{0.76\scriptstyle{\pm 0.31}}$
DroneCircle ${\color[rgb]{.5,.5,.5}0.72\scriptstyle{\pm 0.04}}$ ${\color[rgb]{.5,.5,.5}3.03\scriptstyle{\pm 0.29}}$ $\bm{0.56\scriptstyle{\pm 0.18}}$ $\bm{0.57\scriptstyle{\pm 0.27}}$ ${\color[rgb]{0,0,1}\bm{0.63\scriptstyle{\pm 0.07}}}$ ${\color[rgb]{0,0,1}\bm{0.98\scriptstyle{\pm 0.27}}}$ ${\color[rgb]{.5,.5,.5}0.8\scriptstyle{\pm 0.07}}$ ${\color[rgb]{.5,.5,.5}3.07\scriptstyle{\pm 0.89}}$ ${\color[rgb]{.5,.5,.5}0.78\scriptstyle{\pm 0.04}}$ ${\color[rgb]{.5,.5,.5}3.68\scriptstyle{\pm 0.44}}$ ${\color[rgb]{.5,.5,.5}-0.22\scriptstyle{\pm 0.05}}$ ${\color[rgb]{.5,.5,.5}1.28\scriptstyle{\pm 0.97}}$ ${\color[rgb]{.5,.5,.5}0.26\scriptstyle{\pm 0.03}}$ ${\color[rgb]{.5,.5,.5}1.02\scriptstyle{\pm 0.46}}$ $\bm{0.6\scriptstyle{\pm 0.04}}$ $\bm{0.67\scriptstyle{\pm 0.15}}$
AntCircle ${\color[rgb]{.5,.5,.5}0.58\scriptstyle{\pm 0.19}}$ ${\color[rgb]{.5,.5,.5}4.9\scriptstyle{\pm 3.55}}$ $\bm{0.4\scriptstyle{\pm 0.16}}$ $\bm{0.96\scriptstyle{\pm 2.67}}$ ${\color[rgb]{.5,.5,.5}0.54\scriptstyle{\pm 0.2}}$ ${\color[rgb]{.5,.5,.5}1.78\scriptstyle{\pm 4.33}}$ ${\color[rgb]{.5,.5,.5}0.58\scriptstyle{\pm 0.25}}$ ${\color[rgb]{.5,.5,.5}2.87\scriptstyle{\pm 3.08}}$ ${\color[rgb]{.5,.5,.5}0.65\scriptstyle{\pm 0.2}}$ ${\color[rgb]{.5,.5,.5}5.48\scriptstyle{\pm 3.33}}$ $\bm{0.0\scriptstyle{\pm 0.0}}$ $\bm{0.0\scriptstyle{\pm 0.0}}$ ${\color[rgb]{.5,.5,.5}0.17\scriptstyle{\pm 0.1}}$ ${\color[rgb]{.5,.5,.5}5.04\scriptstyle{\pm 6.74}}$ ${\color[rgb]{0,0,1}\bm{0.47\scriptstyle{\pm 0.04}}}$ ${\color[rgb]{0,0,1}\bm{0.98\scriptstyle{\pm 0.3}}}$
BulletGym Average ${\color[rgb]{.5,.5,.5}0.64\scriptstyle{\pm 0.25}}$ ${\color[rgb]{.5,.5,.5}3.36\scriptstyle{\pm 3.31}}$ $\bm{0.52\scriptstyle{\pm 0.27}}$ $\bm{0.82\scriptstyle{\pm 1.27}}$ ${\color[rgb]{.5,.5,.5}0.68\scriptstyle{\pm 0.19}}$ ${\color[rgb]{.5,.5,.5}1.04\scriptstyle{\pm 1.65}}$ ${\color[rgb]{.5,.5,.5}0.74\scriptstyle{\pm 0.25}}$ ${\color[rgb]{.5,.5,.5}3.11\scriptstyle{\pm 3.55}}$ ${\color[rgb]{.5,.5,.5}0.48\scriptstyle{\pm 0.27}}$ ${\color[rgb]{.5,.5,.5}3.8\scriptstyle{\pm 3.95}}$ ${\color[rgb]{.5,.5,.5}0.33\scriptstyle{\pm 0.29}}$ ${\color[rgb]{.5,.5,.5}1.12\scriptstyle{\pm 1.85}}$ ${\color[rgb]{.5,.5,.5}0.55\scriptstyle{\pm 0.24}}$ ${\color[rgb]{.5,.5,.5}2.55\scriptstyle{\pm 3.62}}$ ${\color[rgb]{0,0,1}\bm{0.61\scriptstyle{\pm 0.19}}}$ ${\color[rgb]{0,0,1}\bm{0.52\scriptstyle{\pm 0.38}}}$
easysparse ${\color[rgb]{.5,.5,.5}0.17\scriptstyle{\pm 0.05}}$ ${\color[rgb]{.5,.5,.5}1.54\scriptstyle{\pm 1.38}}$ $\bm{0.11\scriptstyle{\pm 0.08}}$ $\bm{0.21\scriptstyle{\pm 0.02}}$ $\bm{0.17\scriptstyle{\pm 0.14}}$ $\bm{0.23\scriptstyle{\pm 0.32}}$ ${\color[rgb]{.5,.5,.5}0.78\scriptstyle{\pm 0.0}}$ ${\color[rgb]{.5,.5,.5}5.01\scriptstyle{\pm 0.06}}$ $\bm{0.11\scriptstyle{\pm 0.0}}$ $\bm{0.86\scriptstyle{\pm 0.01}}$ $\bm{-0.06\scriptstyle{\pm 0.0}}$ $\bm{0.07\scriptstyle{\pm 0.02}}$ ${\color[rgb]{.5,.5,.5}0.96\scriptstyle{\pm 0.02}}$ ${\color[rgb]{.5,.5,.5}5.44\scriptstyle{\pm 0.27}}$ ${\color[rgb]{0,0,1}\bm{0.35\scriptstyle{\pm 0.11}}}$ ${\color[rgb]{0,0,1}\bm{0.05\scriptstyle{\pm 0.03}}}$
easymean ${\color[rgb]{.5,.5,.5}0.43\scriptstyle{\pm 0.02}}$ ${\color[rgb]{.5,.5,.5}2.82\scriptstyle{\pm 0.0}}$ $\bm{0.04\scriptstyle{\pm 0.03}}$ $\bm{0.29\scriptstyle{\pm 0.02}}$ ${\color[rgb]{0,0,1}\bm{0.45\scriptstyle{\pm 0.11}}}$ ${\color[rgb]{0,0,1}\bm{0.54\scriptstyle{\pm 0.55}}}$ ${\color[rgb]{.5,.5,.5}0.71\scriptstyle{\pm 0.06}}$ ${\color[rgb]{.5,.5,.5}3.44\scriptstyle{\pm 0.35}}$ $\bm{0.08\scriptstyle{\pm 0.0}}$ $\bm{0.86\scriptstyle{\pm 0.01}}$ $\bm{-0.07\scriptstyle{\pm 0.0}}$ $\bm{0.07\scriptstyle{\pm 0.01}}$ ${\color[rgb]{.5,.5,.5}0.66\scriptstyle{\pm 0.16}}$ ${\color[rgb]{.5,.5,.5}3.97\scriptstyle{\pm 1.47}}$ $\bm{0.29\scriptstyle{\pm 0.04}}$ $\bm{0.05\scriptstyle{\pm 0.03}}$
easydense ${\color[rgb]{.5,.5,.5}0.27\scriptstyle{\pm 0.14}}$ ${\color[rgb]{.5,.5,.5}1.94\scriptstyle{\pm 1.18}}$ $\bm{0.11\scriptstyle{\pm 0.07}}$ $\bm{0.14\scriptstyle{\pm 0.01}}$ ${\color[rgb]{0,0,1}\bm{0.32\scriptstyle{\pm 0.18}}}$ ${\color[rgb]{0,0,1}\bm{0.62\scriptstyle{\pm 0.43}}}$ $\bm{0.26\scriptstyle{\pm 0.0}}$ $\bm{0.47\scriptstyle{\pm 0.01}}$ $\bm{0.02\scriptstyle{\pm 0.05}}$ $\bm{0.41\scriptstyle{\pm 0.22}}$ $\bm{-0.06\scriptstyle{\pm 0.0}}$ $\bm{0.03\scriptstyle{\pm 0.01}}$ ${\color[rgb]{.5,.5,.5}0.5\scriptstyle{\pm 0.1}}$ ${\color[rgb]{.5,.5,.5}2.54\scriptstyle{\pm 0.53}}$ $\bm{0.24\scriptstyle{\pm 0.06}}$ $\bm{0.06\scriptstyle{\pm 0.03}}$
[…]

### v5

@@ Appendix B Extended Results
BallRun ${\color[rgb]{.5,.5,.5}0.6\scriptstyle{\pm 0.1}}$ ${\color[rgb]{.5,.5,.5}5.08\scriptstyle{\pm 0.74}}$ ${\color[rgb]{.5,.5,.5}0.27\scriptstyle{\pm 0.14}}$ ${\color[rgb]{.5,.5,.5}1.46\scriptstyle{\pm 0.39}}$ ${\color[rgb]{.5,.5,.5}0.39\scriptstyle{\pm 0.09}}$ ${\color[rgb]{.5,.5,.5}1.16\scriptstyle{\pm 0.19}}$ ${\color[rgb]{.5,.5,.5}0.76\scriptstyle{\pm 0.01}}$ ${\color[rgb]{.5,.5,.5}3.91\scriptstyle{\pm 0.35}}$ ${\color[rgb]{.5,.5,.5}-0.47\scriptstyle{\pm 0.0}}$ ${\color[rgb]{.5,.5,.5}5.03\scriptstyle{\pm 0.0}}$ ${\color[rgb]{.5,.5,.5}0.22\scriptstyle{\pm 0.0}}$ ${\color[rgb]{.5,.5,.5}1.27\scriptstyle{\pm 0.12}}$ $0.59\scriptstyle{\pm 0.0}$ ${\color[rgb]{.5,.5,.5}3.52\scriptstyle{\pm 0.0}}$ ${\color[rgb]{0,0,1}\bm{0.27\scriptstyle{\pm 0.02}}}$ ${\color[rgb]{0,0,1}\bm{0.49\scriptstyle{\pm 0.13}}}$
CarRun $\bm{0.97\scriptstyle{\pm 0.02}}$ $\bm{0.33\scriptstyle{\pm 0.05}}$ $\bm{0.94\scriptstyle{\pm 0.0}}$ $\bm{0.22\scriptstyle{\pm 0.02}}$ ${\color[rgb]{0,0,1}\bm{0.99\scriptstyle{\pm 0.01}}}$ ${\color[rgb]{0,0,1}\bm{0.65\scriptstyle{\pm 0.31}}}$ $\bm{0.94\scriptstyle{\pm 0.01}}$ $\bm{0.15\scriptstyle{\pm 0.91}}$ ${\color[rgb]{.5,.5,.5}0.68\scriptstyle{\pm 0.01}}$ $7.78\scriptstyle{\pm 0.09}$ ${\color[rgb]{.5,.5,.5}0.95\scriptstyle{\pm 0.01}}$ $1.79\scriptstyle{\pm 0.18}$ $\bm{0.87\scriptstyle{\pm 0.0}}$ $\bm{0.0\scriptstyle{\pm 0.0}}$ $\bm{0.97\scriptstyle{\pm 0.0}}$ $\bm{0.0\scriptstyle{\pm 0.01}}$
DroneRun ${\color[rgb]{.5,.5,.5}0.24\scriptstyle{\pm 0.02}}$ ${\color[rgb]{.5,.5,.5}2.13\scriptstyle{\pm 0.62}}$ $\bm{0.28\scriptstyle{\pm 0.25}}$ $\bm{0.74\scriptstyle{\pm 0.97}}$ ${\color[rgb]{0,0,1}\bm{0.63\scriptstyle{\pm 0.04}}}$ ${\color[rgb]{0,0,1}\bm{0.79\scriptstyle{\pm 0.68}}}$ ${\color[rgb]{.5,.5,.5}0.72\scriptstyle{\pm 0.12}}$ ${\color[rgb]{.5,.5,.5}5.54\scriptstyle{\pm 0.81}}$ ${\color[rgb]{.5,.5,.5}0.42\scriptstyle{\pm 0.1}}$ ${\color[rgb]{.5,.5,.5}2.47\scriptstyle{\pm 0.34}}$ ${\color[rgb]{.5,.5,.5}0.33\scriptstyle{\pm 0.1}}$ ${\color[rgb]{.5,.5,.5}3.52\scriptstyle{\pm 0.58}}$ ${\color[rgb]{.5,.5,.5}0.67\scriptstyle{\pm 0.02}}$ ${\color[rgb]{.5,.5,.5}4.15\scriptstyle{\pm 0.1}}$ $\bm{0.55\scriptstyle{\pm 0.0}}$ $\bm{0.01\scriptstyle{\pm 0.02}}$
AntRun ${\color[rgb]{.5,.5,.5}0.7\scriptstyle{\pm 0.06}}$ ${\color[rgb]{.5,.5,.5}2.93\scriptstyle{\pm 2.4}}$ ${\color[rgb]{.5,.5,.5}0.65\scriptstyle{\pm 0.15}}$ ${\color[rgb]{.5,.5,.5}1.09\scriptstyle{\pm 0.84}}$ ${\color[rgb]{0,0,1}\bm{0.72\scriptstyle{\pm 0.04}}}$ ${\color[rgb]{0,0,1}\bm{0.91\scriptstyle{\pm 0.42}}}$ ${\color[rgb]{.5,.5,.5}0.76\scriptstyle{\pm 0.07}}$ ${\color[rgb]{.5,.5,.5}5.11\scriptstyle{\pm 2.39}}$ $\bm{0.15\scriptstyle{\pm 0.02}}$ $\bm{0.73\scriptstyle{\pm 0.07}}$ $\bm{0.03\scriptstyle{\pm 0.02}}$ $\bm{0.02\scriptstyle{\pm 0.09}}$ $\bm{0.61\scriptstyle{\pm 0.01}}$ $\bm{0.94\scriptstyle{\pm 0.69}}$ $\bm{0.69\scriptstyle{\pm 0.01}}$ $\bm{0.76\scriptstyle{\pm 0.22}}$
BallCircle ${\color[rgb]{.5,.5,.5}0.74\scriptstyle{\pm 0.15}}$ ${\color[rgb]{.5,.5,.5}4.71\scriptstyle{\pm 1.79}}$ $\bm{0.52\scriptstyle{\pm 0.08}}$ $\bm{0.65\scriptstyle{\pm 0.17}}$ ${\color[rgb]{.5,.5,.5}0.77\scriptstyle{\pm 0.06}}$ ${\color[rgb]{.5,.5,.5}1.07\scriptstyle{\pm 0.27}}$ ${\color[rgb]{.5,.5,.5}0.69\scriptstyle{\pm 0.11}}$ ${\color[rgb]{.5,.5,.5}2.36\scriptstyle{\pm 1.04}}$ ${\color[rgb]{.5,.5,.5}0.86\scriptstyle{\pm 0.18}}$ ${\color[rgb]{.5,.5,.5}3.09\scriptstyle{\pm 1.53}}$ ${\color[rgb]{.5,.5,.5}0.64\scriptstyle{\pm 0.01}}$ ${\color[rgb]{.5,.5,.5}0.76\scriptstyle{\pm 0.0}}$ ${\color[rgb]{.5,.5,.5}0.7\scriptstyle{\pm 0.04}}$ ${\color[rgb]{.5,.5,.5}2.61\scriptstyle{\pm 0.79}}$ ${\color[rgb]{0,0,1}\bm{0.69\scriptstyle{\pm 0.06}}}$ ${\color[rgb]{0,0,1}\bm{0.66\scriptstyle{\pm 0.12}}}$
CarCircle ${\color[rgb]{.5,.5,.5}0.58\scriptstyle{\pm 0.25}}$ ${\color[rgb]{.5,.5,.5}3.74\scriptstyle{\pm 2.2}}$ $\bm{0.5\scriptstyle{\pm 0.22}}$ $\bm{0.84\scriptstyle{\pm 0.67}}$ ${\color[rgb]{0,0,1}\bm{0.75\scriptstyle{\pm 0.06}}}$ ${\color[rgb]{0,0,1}\bm{0.95\scriptstyle{\pm 0.61}}}$ ${\color[rgb]{.5,.5,.5}0.63\scriptstyle{\pm 0.19}}$ ${\color[rgb]{.5,.5,.5}1.89\scriptstyle{\pm 1.37}}$ ${\color[rgb]{.5,.5,.5}0.74\scriptstyle{\pm 0.1}}$ ${\color[rgb]{.5,.5,.5}2.18\scriptstyle{\pm 1.33}}$ $\bm{0.71\scriptstyle{\pm 0.02}}$ $\bm{0.33\scriptstyle{\pm 0.0}}$ ${\color[rgb]{.5,.5,.5}0.49\scriptstyle{\pm 0.05}}$ ${\color[rgb]{.5,.5,.5}3.14\scriptstyle{\pm 2.98}}$ $\bm{0.61\scriptstyle{\pm 0.06}}$ $\bm{0.87\scriptstyle{\pm 0.35}}$
DroneCircle ${\color[rgb]{.5,.5,.5}0.72\scriptstyle{\pm 0.04}}$ ${\color[rgb]{.5,.5,.5}3.03\scriptstyle{\pm 0.29}}$ $\bm{0.56\scriptstyle{\pm 0.18}}$ $\bm{0.57\scriptstyle{\pm 0.27}}$ ${\color[rgb]{0,0,1}\bm{0.63\scriptstyle{\pm 0.07}}}$ ${\color[rgb]{0,0,1}\bm{0.98\scriptstyle{\pm 0.27}}}$ ${\color[rgb]{.5,.5,.5}0.8\scriptstyle{\pm 0.07}}$ ${\color[rgb]{.5,.5,.5}3.07\scriptstyle{\pm 0.89}}$ ${\color[rgb]{.5,.5,.5}0.78\scriptstyle{\pm 0.04}}$ ${\color[rgb]{.5,.5,.5}3.68\scriptstyle{\pm 0.44}}$ ${\color[rgb]{.5,.5,.5}-0.22\scriptstyle{\pm 0.05}}$ ${\color[rgb]{.5,.5,.5}1.28\scriptstyle{\pm 0.97}}$ ${\color[rgb]{.5,.5,.5}0.26\scriptstyle{\pm 0.03}}$ ${\color[rgb]{.5,.5,.5}1.02\scriptstyle{\pm 0.46}}$ $\bm{0.6\scriptstyle{\pm 0.05}}$ $\bm{0.66\scriptstyle{\pm 0.15}}$
AntCircle ${\color[rgb]{.5,.5,.5}0.58\scriptstyle{\pm 0.19}}$ ${\color[rgb]{.5,.5,.5}4.9\scriptstyle{\pm 3.55}}$ $\bm{0.4\scriptstyle{\pm 0.16}}$ $\bm{0.96\scriptstyle{\pm 2.67}}$ ${\color[rgb]{.5,.5,.5}0.54\scriptstyle{\pm 0.2}}$ ${\color[rgb]{.5,.5,.5}1.78\scriptstyle{\pm 4.33}}$ ${\color[rgb]{.5,.5,.5}0.58\scriptstyle{\pm 0.25}}$ ${\color[rgb]{.5,.5,.5}2.87\scriptstyle{\pm 3.08}}$ ${\color[rgb]{.5,.5,.5}0.65\scriptstyle{\pm 0.2}}$ ${\color[rgb]{.5,.5,.5}5.48\scriptstyle{\pm 3.33}}$ $\bm{0.0\scriptstyle{\pm 0.0}}$ $\bm{0.0\scriptstyle{\pm 0.0}}$ ${\color[rgb]{.5,.5,.5}0.17\scriptstyle{\pm 0.1}}$ ${\color[rgb]{.5,.5,.5}5.04\scriptstyle{\pm 6.74}}$ ${\color[rgb]{0,0,1}\bm{0.49\scriptstyle{\pm 0.03}}}$ ${\color[rgb]{0,0,1}\bm{0.91\scriptstyle{\pm 0.39}}}$
BulletGym Average ${\color[rgb]{.5,.5,.5}0.64\scriptstyle{\pm 0.25}}$ ${\color[rgb]{.5,.5,.5}3.36\scriptstyle{\pm 3.31}}$ $\bm{0.52\scriptstyle{\pm 0.27}}$ $\bm{0.82\scriptstyle{\pm 1.27}}$ ${\color[rgb]{.5,.5,.5}0.68\scriptstyle{\pm 0.19}}$ ${\color[rgb]{.5,.5,.5}1.04\scriptstyle{\pm 1.65}}$ ${\color[rgb]{.5,.5,.5}0.74\scriptstyle{\pm 0.25}}$ ${\color[rgb]{.5,.5,.5}3.11\scriptstyle{\pm 3.55}}$ ${\color[rgb]{.5,.5,.5}0.48\scriptstyle{\pm 0.27}}$ ${\color[rgb]{.5,.5,.5}3.8\scriptstyle{\pm 3.95}}$ ${\color[rgb]{.5,.5,.5}0.33\scriptstyle{\pm 0.29}}$ ${\color[rgb]{.5,.5,.5}1.12\scriptstyle{\pm 1.85}}$ ${\color[rgb]{.5,.5,.5}0.55\scriptstyle{\pm 0.24}}$ ${\color[rgb]{.5,.5,.5}2.55\scriptstyle{\pm 3.62}}$ ${\color[rgb]{0,0,1}\bm{0.61\scriptstyle{\pm 0.19}}}$ ${\color[rgb]{0,0,1}\bm{0.55\scriptstyle{\pm 0.4}}}$
easysparse ${\color[rgb]{.5,.5,.5}0.17\scriptstyle{\pm 0.05}}$ ${\color[rgb]{.5,.5,.5}1.54\scriptstyle{\pm 1.38}}$ $\bm{0.11\scriptstyle{\pm 0.08}}$ $\bm{0.21\scriptstyle{\pm 0.02}}$ $\bm{0.17\scriptstyle{\pm 0.14}}$ $\bm{0.23\scriptstyle{\pm 0.32}}$ ${\color[rgb]{.5,.5,.5}0.78\scriptstyle{\pm 0.0}}$ ${\color[rgb]{.5,.5,.5}5.01\scriptstyle{\pm 0.06}}$ $\bm{0.11\scriptstyle{\pm 0.0}}$ $\bm{0.86\scriptstyle{\pm 0.01}}$ $\bm{-0.06\scriptstyle{\pm 0.0}}$ $\bm{0.07\scriptstyle{\pm 0.02}}$ ${\color[rgb]{.5,.5,.5}0.96\scriptstyle{\pm 0.02}}$ ${\color[rgb]{.5,.5,.5}5.44\scriptstyle{\pm 0.27}}$ ${\color[rgb]{0,0,1}\bm{0.45\scriptstyle{\pm 0.22}}}$ ${\color[rgb]{0,0,1}\bm{0.52\scriptstyle{\pm 0.34}}}$
easymean ${\color[rgb]{.5,.5,.5}0.43\scriptstyle{\pm 0.02}}$ ${\color[rgb]{.5,.5,.5}2.82\scriptstyle{\pm 0.0}}$ $\bm{0.04\scriptstyle{\pm 0.03}}$ $\bm{0.29\scriptstyle{\pm 0.02}}$ ${\color[rgb]{0,0,1}\bm{0.45\scriptstyle{\pm 0.11}}}$ ${\color[rgb]{0,0,1}\bm{0.54\scriptstyle{\pm 0.55}}}$ ${\color[rgb]{.5,.5,.5}0.71\scriptstyle{\pm 0.06}}$ ${\color[rgb]{.5,.5,.5}3.44\scriptstyle{\pm 0.35}}$ $\bm{0.08\scriptstyle{\pm 0.0}}$ $\bm{0.86\scriptstyle{\pm 0.01}}$ $\bm{-0.07\scriptstyle{\pm 0.0}}$ $\bm{0.07\scriptstyle{\pm 0.01}}$ ${\color[rgb]{.5,.5,.5}0.66\scriptstyle{\pm 0.16}}$ ${\color[rgb]{.5,.5,.5}3.97\scriptstyle{\pm 1.47}}$ $\bm{0.4\scriptstyle{\pm 0.21}}$ $\bm{0.47\scriptstyle{\pm 0.36}}$
easydense ${\color[rgb]{.5,.5,.5}0.27\scriptstyle{\pm 0.14}}$ ${\color[rgb]{.5,.5,.5}1.94\scriptstyle{\pm 1.18}}$ $\bm{0.11\scriptstyle{\pm 0.07}}$ $\bm{0.14\scriptstyle{\pm 0.01}}$ $\bm{0.32\scriptstyle{\pm 0.18}}$ $\bm{0.62\scriptstyle{\pm 0.43}}$ $\bm{0.26\scriptstyle{\pm 0.0}}$ $\bm{0.47\scriptstyle{\pm 0.01}}$ $\bm{0.02\scriptstyle{\pm 0.05}}$ $\bm{0.41\scriptstyle{\pm 0.22}}$ $\bm{-0.06\scriptstyle{\pm 0.0}}$ $\bm{0.03\scriptstyle{\pm 0.01}}$ ${\color[rgb]{.5,.5,.5}0.5\scriptstyle{\pm 0.1}}$ ${\color[rgb]{.5,.5,.5}2.54\scriptstyle{\pm 0.53}}$ ${\color[rgb]{0,0,1}\bm{0.37\scriptstyle{\pm 0.08}}}$ ${\color[rgb]{0,0,1}\bm{0.41\scriptstyle{\pm 0.21}}}$
[…]